# Synesthesia of Machines (SoM)-Enhanced ISAC Precoding for Vehicular Networks with Double Dynamics

### v1

@@ V-F Network Update
TABLE III: Hyper-Parameters in proposed scheme
@@
Batch size $N_{b}$ 64
Learning rate $\kappa$ 0.02
discount factor $\gamma$ 0.6
@@
Initial $\lambda$ 0.5
@@
(a) Reward of different schemes.
(b) Reward from different inputs.
Figure 6: Cumulative reward comparison among different schemes.
Then we set the unified loss function for PSAC as $Z^{(k)}(\phi)=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}Z_{i}^{(k)}$, where
$\phi^{(k)}$ $=\phi^{(k-1)}-\kappa\nabla_{\phi}Z^{(k)}(\phi)$ (48)
(a) Averaged CRLB of different schemes.
(b) Averaged SE of different schemes.
Figure 7: ISAC performance comparison among different schemes.
@@ VI Simulation Results
In this section, experimental results are presented to evaluate the performance of the proposed ISAC precoding schemes in doubly-dynamic scenarios. The system setup for environment generation and the hyper-parameters setup for proposed DRL-aided scheme in simulations can be found in Table II and Table III, respectively. The methods to be evaluated includes:
@@ VI-A Cumulative Reward Versus Episodes in the Training
Firstly, we verify BS’s achievable cumulative reward within one frame in the training process of the proposed schemes in Fig. 6. We set SNR as 0 dB and $U=8$. As the number of episodes increases, the cumulative reward of the proposed algorithm gradually rises. After convergence, the cumulative reward of the DRL-PC can surpass the optimization-based method with perfect state information and MP-based benchmarks. This reflects that the proposed DRL-aided method can alleviate the performance loss from the relaxation in optimization-based algorithms and mitigate the reliance on instantaneous state information in double dynamics.
Compared to the training of the classic actor-critic frameworks, the adopted PSAC architecture exhibits much faster convergence speed, illustrating its advantage in accelerating training in complicated state space and action space. Additionally, DRL-AU scheme exhibits a faster convergence compared to DRL-UU scheme, reflecting its higher updating efficiency.
Even when relying solely on historical position estimates in DRL-PO version, the proposed scheme still achieves high reward, keeping a narrow gap between DRL-PC version. However, merely initial CSI as input in DRL-CO is not enough for BS in the training, leading to degraded and unstable cumulative reward.
@@ VI-B ISAC Performance in Different Schemes
With respect to communications, we compare the averaged SE of different schemes. Although BS has no access to the instantaneous CSI at each subframe, it can adjust the precoding matrix through interaction with the environment to gradually keep up with the changes of the channels. The proposed algorithm consistently outperforms the Opt-based method by a performance gain of almost 2.5 dB.
@@ VI-E Impact of Environmental Double Dynamics
Figure 14: ISAC performance boundary of different methods.
Figure 15: Running time comparisons.
@@ VI-G Complexity Analysis and Comparison
The computational complexity of different schemes are compared in Table. IV. The required number of floating point operations (FLOPs) is used as the metric for complexity. In the online deployment of the proposed DRL-aided scheme, only the trained actor is activated for updating $\bm{F}_{\text{RF}}$ and $\bm{F}_{\text{BB}}$ via forward pass. Denote $L_{\text{c}}$ as the maximal number of convolutional layers, $N_{\text{f}}$ as the maximal number of input and output feature maps, and $N_{\text{k}}$ as the maximal side length of the filters in convolutional layers. $N_{\text{W}}$ and $N_{\text{L}}$ as the maximal number of neurons in the hidden layers and the total number of hidden layers in fully-connected layers. Then the complexity in the forward pass of the actor network is $\mathcal{O}(L_{\text{c}}(UG_{\text{t}}N_{\text{d}}+N_{\text{x}}N_{\text{y}})N_{\text{k}}^{2}N_{\text{f}}^{2}+N_{\text{L}}^{2}N_{\text{W}})$, which is linear to $U$ and independent with $N_{\text{t}}$, $M$. Based on the action, the precoding is updated with the complexity $\mathcal{O}(N_{t}+M)$ for DRL-UU and $\mathcal{O}(UM)$ for DRL-AU.
We further compared the averaged running times within one frame of different schemes in Fig. 15. The proposed DRL-aided algorithm demonstrates minimal complexity and considerably shorter running times compared to other benchmarks, highlighting its advantage in reducing complexity. Additionally, the DRL-AU scheme boasts a shorter runtime in comparison to the DRL-UU scheme, demonstrating its higher precoding update efficiency.
TABLE IV: Complexity comparison
@@ VII Conclusion
This paper investigated the ISAC hybrid precoding for scenarios with time-varying communication and sensing channels. We proposed an optimization-based method for wideband multi-user ISAC systems with full environmental knowledge and a SoM-enhanced DRL-based precoding approach for cases with imperfect state information. By leveraging multi-modalities, including CSI and GPS, our method adapts to dynamic changes effectively. The enhanced training algorithm is able to navigate complex action spaces and adjust to varying user numbers. The devised hybrid precoding update scheme spans both user and antenna dimensions, leading to improved ISAC performance in doubly-dynamic scenarios while reducing reliance on real-time information and computational burdens. The multifaceted advantages were validated by extensive simulations.

### v2

@@ V-G Convergence Analysis
TABLE IV: Hyper-parameters in model training
@@
Discount factor $\gamma$ 0.6
Exploration rate $\xi$ 0.2
Batch size $N_{\text{b}}$ 64
Learning rate $\kappa$ 0.02
@@
Initial critic reliability $\lambda$ 0.5
@@ VI Simulation Results
In this section, experimental results are presented to evaluate the performance of the proposed ISAC precoding schemes in doubly-dynamic scenarios.
We first construct a joint GPS-CSI dataset to simulate doubly dynamic vehicular network scenarios, which serves as the platform for training and testing the algorithms.
  • Creation of the scenario: We utilize the Simulation of Urban MObility (SUMO) software to simulate the dynamic trajectories of both moving vehicles and the low-altitude UAV at an intersection [45], where the road width is 30 m. The BS covers a lane length of 180 meters horizontally and 80 meters vertically. The vehicle trajectories include straight motion, turning and lane-changing, while the UAV trajectories encompass horizontal straight-line flight, vertical lift and hovering.
  • Construction of prior positioning information: Once the scene is constructed, the coordinates of the vehicles and the UAV within the intersection are exported. The GPS feedbacks from the users are generated by adding noise to their coordinates, simulating GPS errors in real-world scenarios.
  • Generation of channel data: After importing the scene into Wireless InSite® software [46], communication and sensing channels for each symbol duration are generated using ray-tracing, based on the channel model outlined in Section II. This channel information serves as the foundation for calculating the loss and training the model.
The system setup for environment generation and the hyper-parameters setup for proposed DRL-aided scheme in simulations can be found in Table III and Table IV, respectively. The methods to be evaluated includes:
@@
(a) Reward of different schemes.
(b) Reward from different inputs.
Figure 6: Cumulative reward comparison among different schemes.
@@ VI-A Cumulative Reward Versus Episodes in the Training
Firstly, we verify BS’s achievable cumulative reward within one frame in the training process of the proposed schemes in Fig. 6. We set SNR as 0 dB and $U=8$. As the number of episodes increases, the cumulative reward of the proposed algorithm gradually rises. After convergence, the cumulative reward of the DRL-PC can surpass the optimization-based method with perfect state information and MP-based benchmarks. This reflects that the proposed DRL-aided method can alleviate the performance loss from the relaxation in optimization-based algorithms and mitigate the reliance on instantaneous state information in double dynamics. Even when relying solely on historical position estimates in DRL-PO version, the proposed scheme still achieves high reward, keeping a narrow gap between DRL-PC. However, merely initial CSI as input in DRL-CO is not enough for BS in the training, leading to degraded and unstable cumulative reward.
Compared to the training of the classic actor-critic frameworks, the adopted PSAC architecture exhibits much faster convergence speed, illustrating its advantage in accelerating training in complicated state space and action space. Additionally, DRL-AU scheme exhibits a faster convergence compared to DRL-UU scheme, reflecting its higher updating efficiency. Therefore, the DRL-UU is better suited for environments with more stable conditions, while the DRL-AU is more appropriate for dynamic scenarios involving rapid changes in the distribution of targets and users.
We also compare the impact of hyper-parameters on the convergence speed and the achievable rewards in the training of DRL-AU in Table. V and Table. VI. Both excessively high and low discount factors $\gamma$ lead to reduced performance after convergence. A higher $\xi$ results in a slower convergence, while a lower $\xi$ accelerates convergence but causes the model to reside at a local optimum, thereby reducing the achievable cumulative reward. Therefore, appropriately selecting hyper-parameters is crucial for improving both model training efficiency and performance.
TABLE V: Impact of $\gamma$ on the training
$\gamma$ 0 0.2 0.4 0.6 0.8 1.0
Converging epochs 1810 1840 1870 1880 1880 1910
Cumulative reward 84.69 88.52 90.77 91.56 90.34 87.21
TABLE VI: Impact of $\xi$ on the training
$\xi$ 0 0.2 0.4 0.6 0.8 1.0
Converging epochs 1710 1880 2420 3510 4180 -
Cumulative reward 86.04 91.56 91.57 91.57 91.57 -
(a) Averaged CRLB of different schemes.
(b) Averaged SE of different schemes.
Figure 7: ISAC performance comparison among different schemes.
@@ VI-B ISAC Performance in Different Schemes
With respect to communications, we compare the averaged SE of different schemes. Despite lacking the instantaneous CSI at each subframe, the BS can adjust the precoding matrix through interaction with the environment to gradually keep up with the changes of the channels. The proposed algorithm consistently outperforms the Opt-based method by a performance gain of almost 2.5 dB.
Moreover, the ISAC performance of the DRL-aided scheme closely approaches the ideal case when perfect channel prior information is used as input rather than the initial CSI. This suggests that the proposed scheme does not purely rely on perfect prior information. Additionally, the performance in this case surpasses that of the Opt-based approach, providing a more equitable comparison that highlights the superiority of the DRL-aided scheme.
@@ VI-G Complexity Analysis and Comparison
The computational complexity of different schemes are compared in Table. VII. The required number of floating point operations (FLOPs) is used as the metric for complexity. In the online deployment of the proposed DRL-aided scheme, only the trained actor is activated for updating $\bm{F}_{\text{RF}}$ and $\bm{F}_{\text{BB}}$ via forward pass. Denote $L_{\text{c}}$ as the maximal number of convolutional layers, $N_{\text{f}}$ as the maximal number of input and output feature maps, and $N_{\text{k}}$ as the maximal side length of the filters in convolutional layers. $N_{\text{W}}$ and $N_{\text{L}}$ as the maximal number of neurons in the hidden layers and the total number of hidden layers in fully-connected layers. Then the complexity in the forward pass of the actor network is $\mathcal{O}(L_{\text{c}}(UG_{\text{t}}N_{\text{d}}+N_{\text{x}}N_{\text{y}})N_{\text{k}}^{2}N_{\text{f}}^{2}+N_{\text{L}}^{2}N_{\text{W}})$, which is linear to $U$ and independent with $N_{\text{t}}$, $M$. Based on the action, the precoding is updated with the complexity $\mathcal{O}(N_{t}+M)$ for DRL-UU and $\mathcal{O}(UM)$ for DRL-AU. Hence, the DRL-UU demonstrates lower complexity in systems with a large number of users, whereas the DRL-AU exhibits lower complexity in systems with large antenna arrays.
We further compared the averaged running times within one frame of different schemes in Fig. 15. The proposed DRL-aided algorithm demonstrates minimal complexity and considerably shorter running times compared to other benchmarks, highlighting its advantage in reducing complexity. Additionally, the DRL-AU scheme boasts a shorter runtime in comparison to the DRL-UU scheme.
Figure 14: ISAC performance boundary of different methods.
Figure 15: Running time comparisons.
TABLE VII: Complexity comparison
@@ VII Conclusion
This paper investigated the ISAC hybrid precoding for scenarios with time-varying communication and sensing channels. We proposed an optimization-based method for wideband multi-user ISAC systems with full environmental knowledge and a SoM-enhanced DRL-based precoding approach for the cases with imperfect state information. By leveraging different modalities, including CSI and GPS, our method adapts to dynamic changes effectively. The enhanced training algorithm is able to navigate complex action spaces and adjust to varying user numbers. The devised hybrid precoding update scheme spans both the user and the antenna dimensions, leading to improved ISAC performance in doubly-dynamic scenarios while reducing reliance on real-time information and computational burdens. The multifaceted advantages were validated by extensive simulations.
Appendix A Proof of the proposition 1
Denote $\widehat{\bm{W}}_{m,u}$ as the global optimum to the relaxed version from (15). The precoding vectors are constructed as
$\widetilde{\bm{f}}_{m,u}=(\bm{h}_{m,m,u}^{\text{H}}\widehat{\bm{W}}_{m,u}\bm{h}_{m,m,u})^{-\frac{1}{2}}\widehat{\bm{W}}_{m,u}\bm{h}_{m,m,u},$ (50)
for $m\in[1,M]$ and $u\in[1,U]$. Then the covariance are $\widetilde{\bm{W}}_{m,u}=\widetilde{\bm{f}}_{m,u}\widetilde{\bm{f}}_{m,u}^{\text{H}}$. It is obvious that $\widetilde{\bm{W}}_{m,u}$ is positive semi-definite and rank-1. Since the objective function in (15) is not directly influenced by $\widetilde{\bm{W}}_{m,u}$, we only need to verify that $\widetilde{\bm{W}}_{m,u}$ is a feasible solution to (15).
Firstly, it can be derived that
$\bm{h}^{\text{H}}_{m,m,u}\widetilde{\bm{W}}_{m,u}\bm{h}_{m,m,u}=$ $\bm{h}^{\text{H}}_{m,m,u}\widetilde{\bm{f}}_{m,u}\widetilde{\bm{f}}^{\text{H}}_{m,u}\bm{h}_{m,m,u}$
$=$ $\bm{h}^{\text{H}}_{m,m,u}\widehat{\bm{W}}_{m,u}\bm{h}_{m,m,u},$ (51)
while $\bm{h}^{\text{H}}_{m,k,u}\widetilde{\bm{W}}_{k,i}\bm{h}_{m,k,u}\leqslant\bm{h}^{\text{H}}_{m,k,i}\widetilde{\bm{W}}_{k,i}\bm{h}_{m,k,i}$ when $k\neq m$ or $i\neq u$. Thus $\text{Tr}(\bm{Q}_{m,k,u}\widetilde{\bm{W}}_{k,i})\leq\text{Tr}(\bm{Q}_{m,m,u}\widehat{\bm{W}}_{m,u})$, taking equal if and only if $k=m$ and $i=u$. Substituting these to (15d) gives rise to
$\text{Tr}(\bm{Q}_{\text{m,m,u}}\!\widetilde{\bm{W}}_{\text{m,u}})\!\!-\!\tau\!\!\!\sum_{i=1,i\neq u}^{U}\!\!\text{Tr}(\bm{Q}_{\text{m,m,u}}\widetilde{\bm{W}}_{\text{m,i}})\!\!-\!\tau\!\!\!\sum_{k=1,k\neq m}\!\!\!\!\bm{Q}_{\text{m,k,u}}\widetilde{\bm{W}}_{\text{k,u}}$
$\geqslant\!\!\text{Tr}(\!\bm{Q}_{\text{m,m,u}}\!\widehat{\bm{W}}_{\text{m,u}})\!\!-\!\!\tau\!\!\!\sum_{i=1,i\neq u}^{U}\!\!\!\text{Tr}(\bm{Q}_{\text{m,m,u}}\!\widehat{\bm{W}}_{\text{m,i}})\!\!-\!\tau\!\!\!\sum_{k=1,k\neq m}\!\!\!\!\bm{Q}_{\text{m,k,u}}\!\widehat{\bm{W}}_{\text{k,u}}$
$\geqslant\tau\sigma_{\text{c}}^{2},\vspace{-0.2cm}$ (52)
namely, SE-related constraint (15d) holds for $\widetilde{\bm{W}}_{m,u}$.
Subsequently, we prove that the CRB-related constraint also holds for $\widetilde{\bm{W}}_{m,u}$. Notice that $\text{Tr}(\dot{\bm{A}}^{\text{H}}\dot{\bm{A}}\widetilde{\bm{W}}_{m,u})=\text{Tr}(\dot{\bm{A}}^{\text{H}}\dot{\bm{A}}(\bm{h}_{m,m,u}^{\text{H}}\widehat{\bm{W}}_{m,u}\bm{h}_{m,m,u})^{-1}\widehat{\bm{W}}_{m,u}\bm{h}_{m,m,u}\bm{h}_{m,m,u}^{\text{H}}\widehat{\bm{W}}_{m,u})=\text{Tr}(\dot{\bm{A}}^{\text{H}}\dot{\bm{A}}\widehat{\bm{W}}_{m,u})$, it can be derived that
$\left[\text{Tr}(\dot{\bm{A}}^{\text{H}}\dot{\bm{A}}\widetilde{\bm{W}}_{m,u})\!-\!\zeta\right]\text{Tr}(\bm{A}^{\text{H}}\bm{A}\widetilde{\bm{W}}_{m,u})\!-\!|\text{Tr}(\dot{\bm{A}}^{\text{H}}\bm{A}\widetilde{\bm{W}}_{m,u})|^{2}$
$=$ $\left[\text{Tr}(\dot{\bm{A}}^{\text{H}}\dot{\bm{A}}\widehat{\bm{W}}_{m,u})-\zeta\right]\text{Tr}(\bm{A}^{\text{H}}\bm{A}\widetilde{\bm{W}}_{m,u})\geqslant 0,$ (53)
given the orthogonality between column vectors in $\bm{A}$ and $\dot{\bm{A}}$. By summing (53) over $u$, it can be proved that the constraint (15c) holds.
Finally, we will examine the power constraint as follows:
$\sum_{m=1}^{M}\text{Tr}(\tilde{\bm{W}}_{m})=\sum_{m=1}^{M}\sum_{u=1}^{U}\text{Tr}(\tilde{\bm{W}}_{m,u})$
$=$ $\!\!\!\sum_{m=1}^{M}\!\!\sum_{u=1}^{U}\!(\bm{h}_{m,m,u}^{\text{H}}\!\!\hat{\bm{W}}_{m,u}\bm{h}_{m,m,u}\!)^{-1}\text{Tr}(\bm{h}_{m,m,u}^{\text{H}}\!\!\hat{\bm{W}}_{m,u}^{\text{H}}\!\!\hat{\bm{W}}_{m,u}\bm{h}_{m,m,u}\!)$
$\leqslant$ $\!\!\!\sum_{m=1}^{M}\!\sum_{u=1}^{U}\!\text{Tr}(\bm{h}_{m,m,u}\bm{h}_{m,m,u}^{\text{H}}\!\!\hat{\bm{W}}_{m,u})^{-1}\!\text{Tr}(\!\bm{h}_{m,m,u}\bm{h}_{m,m,u}^{\text{H}}\!\!\hat{\bm{W}}_{m,u})$
$\text{Tr}(\hat{\bm{W}}_{m,u})=\text{Tr}(\hat{\bm{W}}_{m,u})\leqslant P_{\text{t}},\vspace{-0.2cm}$ (54)
according to the operator norm inequality. Hence the power constraint (15e) also holds.
With the derivation above, it is verified that $\tilde{\bm{W}}_{m,u}$ is a feasible solution, and furthermore, it is also a global optimum to (15), thereby completing the proof.
Appendix B Proof of the Proposition 2
Under the assumptions in [44], the policy gradient $\nabla J(\mu)$ is Lipschitz continuous, laying the foundation for the subsequent finite-sample analysis.
First, we characterize the propagation of the dynamics of critic’s tracking error by coupling it with the actor’s updates. Using the time-difference learning property, the update of the critic network’s parameter $\phi_{\text{C}}$ follows
$\mathbb{E}\|\phi_{\text{C}}^{(t+1)}\!-\!\phi_{\text{C}|\mu}^{(t)\star}\|^{2}\!\leqslant\!(1\!-\!\alpha_{1}\kappa_{\text{C}})\mathbb{E}\|\phi_{\text{C}}^{(t)}\!-\!\phi_{\text{C}|\mu}^{(t)\star}\|^{2}+\!\frac{\alpha_{2}\kappa_{\text{C}}^{2}}{N_{\text{b}}},$ (55)
with $\alpha_{1}$, $\alpha_{2}$ being constants. Taking into account the actor’s update renders
$\mathbb{E}$ $\|\phi_{\text{C}}^{(t+1)}-\phi_{\text{C}|\mu}^{(t+1)\star}\|^{2}\leqslant\frac{\alpha_{3}}{\kappa_{\text{C}}}\mathbb{E}\|\phi_{\text{A}}^{(t+1)}-\phi_{\text{A}}^{(t)}\|^{2}$ (56)
$+(1-\frac{\alpha_{1}\kappa_{\text{C}}}{2})\mathbb{E}\|\phi_{\text{C}}^{(t)}-\phi_{\text{C}|\mu}^{(t+1)\star}\|^{2}+\frac{2\alpha_{2}\kappa_{\text{C}}^{2}}{N_{\text{b}}},$
where $\phi_{\text{C}|\mu}^{(t)\star}$ is the global optimum of TD for a given fixed policy for a given $\mu$, and $\alpha_{3}$ is another constant.
Subsequently, the cumulative tracking error can be bounded by considering the dynamics of the tracking error. From the update rule of the actor parameter $\phi_{\text{A}}$, we have
$\phi_{\text{A}}^{(t+1)}-\phi_{\text{A}}^{(t)}$ $=\frac{\kappa_{\text{A}}}{N_{\text{b}}}\sum_{n=1}^{N_{\text{b}}}\nabla_{\phi_{\text{A}}}\mu^{(t)}(\bm{o}_{t,n})\nabla_{\phi_{\text{A}}}\mu^{(t)}(\bm{o}_{t,n})^{\text{T}}\phi_{\text{C}}^{(t)}$
$:=\alpha_{\mu}h_{\mu}(\phi_{\text{C}}^{(t)}).$ (57)
By exploiting the property of the Fisher information of the defined policy, the estimation bias is characterized as
$\vspace{-0.2cm}\mathbb{E}\|h_{\mu}(\phi^{(t)}_{\text{C}})\!-\!\nabla J^{(t)}(\mu)\|^{2}\!\leqslant\!\alpha_{4}\mathbb{E}\|\phi_{\text{C}}^{(t)}\!-\!\phi_{\text{C}|\mu}^{(t)\star}\|^{2}\!+\!\alpha_{4}\kappa^{2}\!+\!\frac{\alpha_{5}}{N_{\text{b}}},$ (58)
with $\alpha_{4}$, $\alpha_{5}$ being constants. Then the cumulative dynamic tracking error is expressed as
$\sum_{t=0}^{T-1}$ $\mathbb{E}\left\lVert\phi_{\text{C}}^{(t)}-\phi_{\text{C}|\mu}^{(t)\star}\right\rVert^{2}\leqslant\frac{4\left\lVert\phi_{\text{C}}^{(0)}-\phi^{\star}_{\text{C}|\mu^{(0)}}\right\rVert^{2}}{\alpha_{1}\kappa_{\text{C}}}$
$+\left[\frac{2\alpha_{2}\kappa_{\text{C}}^{2}}{N_{\text{b}}}+\frac{2\alpha_{3}\kappa_{\text{A}}^{2}}{\kappa_{\text{C}}}\left(\alpha_{4}\kappa^{2}+\frac{\alpha_{5}}{N_{\text{b}}}\right)\right]\frac{4T}{\alpha_{1}\kappa_{\text{C}}}$
$+\frac{\alpha_{6}\kappa_{\text{A}}^{2}}{\kappa_{\text{C}}^{2}}\sum_{t=0}^{T-1}\mathbb{E}\left\lVert\nabla J(\mu^{(t)})\right\rVert^{2},$ (59)
which is related to the convergence speed of the actor’s update towards the stationary point, with $\alpha_{6}$ being another constant.
Ultimately we establish the overall convergence to a stationary policy by mitigating the cumulative tracking error through the actors’ update process. By examining the relationship between the progress of the loss function and the tracking error, we can bound the cumulative policy gradient by
$\frac{\kappa_{\text{A}}}{4}\sum_{t=0}^{T-1}$ $\mathbb{E}\left\lVert\nabla J(\mu^{(t)})\right\rVert^{2}\leqslant\frac{3\alpha_{4}\kappa_{\text{A}}}{4}\sum_{t=0}^{T-1}\mathbb{E}\left\lVert\phi_{\text{C}}^{(t)}-\phi_{\text{C}|\mu}^{(t)\star}\right\rVert^{2}$
$+\frac{R_{\text{max}}}{1\!-\!\gamma}\!+\!\frac{3\kappa_{\text{A}}}{4}\left(\alpha_{4}\kappa^{2}+\frac{\alpha_{5}}{N_{\text{b}}}\right)T.$ (60)
Utilizing the relationship between the cumulative tracking error and the cumulative policy gradient, the tracking error can be ultimately canceled by the actor’s positive progress towards a stationary point:
$\left(\frac{\kappa_{\text{A}}}{4}-\frac{3\alpha_{6}\kappa_{\text{A}}^{3}}{4\kappa_{\text{C}}^{2}}\right)\sum_{t=0}^{T-1}\mathbb{E}\left\lVert\nabla J(\mu^{(t)})\right\rVert^{2}$
$\leqslant\frac{R_{\text{max}}}{1-\gamma}+\frac{3\alpha_{4}\kappa_{\text{A}}}{\alpha_{1}\kappa_{\text{C}}}\left\lVert\phi_{\text{C}}^{(0)}-\phi^{\star}_{\text{C}|\mu^{(0)}}\right\rVert^{2}$
$+\left[\frac{2\alpha_{2}\kappa_{\text{C}}^{2}}{N_{\text{b}}}+\frac{2\alpha_{3}\kappa_{\text{A}}^{2}}{\kappa_{\text{C}}}\left(\alpha_{4}\kappa^{2}+\frac{\alpha_{5}}{N_{\text{b}}}\right)\right]\frac{3\alpha_{4}\kappa_{\text{A}}}{\alpha_{1}\kappa_{\text{C}}}T$
$+\frac{3\kappa_{\text{A}}}{4}\left(\alpha_{4}\kappa^{2}+\frac{\alpha_{5}}{N_{\text{b}}}\right)T.$ (61)
Finally, by rearranging these terms, Eq. (49) in Proposition 2 can be obtained.